OpenStudy (anonymous):

How do I factor x^2-3x-1?

myininaya (myininaya):

not factorable over the integers

OpenStudy (anonymous):

hey myininaya is right

OpenStudy (anonymous):

so then how do I do x^2-3x-1/4x^2-9?

OpenStudy (anonymous):

you can always complete the square

OpenStudy (anonymous):

Because my original question from the book is: 2x^2+3x+1 x^2-3x-1 ____________ - ________________ 4x^2-9 -(4x^2-9)

OpenStudy (anonymous):

I did: (2x+1)(x+1) ____________ + ______ 4x^2-9 4x^2-9

OpenStudy (anonymous):

...

myininaya (myininaya):

but it can still be factored \[b=(-\frac{3}{2}+z)+(-\frac{3}{2}-z)\] \[a*c=(\frac{-3}{2}+z)(\frac{-3}{2}-z)=-1\] so what is z \[\frac{9}{4}-z^2=-1 =>\frac{9}{4}+1=z^2=>z^2=\frac{13}{4}=>z=\pm \frac{\sqrt{13}}{2}\] so we have \[x^2+(\frac{-3}{2}+ \frac{\sqrt{13}}{2})x+(\frac{-3}{2}- \frac{\sqrt{13}}{2})x-1\] \[x(x+\frac{-3 + \sqrt{13}}{2})+\frac{-3- \sqrt{13}}{2}(x-1 \cdot \frac{2}{-3- \sqrt{13}})\] \[x(x+\frac{-3 + \sqrt{13}}{2})+\frac{-3- \sqrt{13}}{2}(x-\frac{2}{-3- \sqrt{13}} \cdot \frac{-3 + \sqrt{13}}{-3+ \sqrt{13}})\] \[x(x+\frac{-3+ \sqrt{13}}{2})+\frac{-3 - \sqrt{13}}{2}(x-\frac{2(-3 + \sqrt{13})}{9-13})\] \[x(x+\frac{-3+ \sqrt{13}}{2})+\frac{-3 - \sqrt{13}}{2}(x-\frac{2(-3 + \sqrt{13})}{-4})\] \[x(x+\frac{-3 + \sqrt{13}}{2})+\frac{-3- \sqrt{13}}{2}(x+\frac{-3+ \sqrt{13}}{2})\] \[(x+\frac{-3+\sqrt{13}}{2})(x+\frac{-3-\sqrt{13}}{2})\]

myininaya (myininaya):

2x^2+3x+1 x^2-3x-1 ____________ - ________________ 4x^2-9 -(4x^2-9) \[\frac{2x^2+3x+1}{4x^2-9}+\frac{x^2-3x-1}{4x^2-9}\]

myininaya (myininaya):

you have same denominator add the numerators

OpenStudy (anonymous):

ohh, okay.